1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

The Banshee (An Approach)

Der Werwolf

El Hechicero

tradukita de Max Knight tradukita de Hans Eiseneck

One night, a banshee slunk away from mate and child, and in the

gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

Ein Werwolf eines Nachts ent-

wich von Weib und Kind und sich be-

gab an eines Dorfschullehrers Grab und bat ihn: �Bitte, beuge

mich!“ Un hechicero, con esposa vino un día a la fosa de su maestro, y al profe-

sor pidió: “¿Me cuentas? ¡Por

favor!”

The village teacher climbed up

straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

Der Dorfschulmeister stieg hin-

auf auf seines Blechschilds Messing-

knauf und sprach zum Wolf, der seine

Pfoten geduldig kreuzte vor dem To-

ten:

El profesor sin vacilar subió y comenzó a hablar. Desde allá de su letrero se dirigió al hechicero:

“The banSHEE, in the subject’s

place; the banHERS, the possessive case. The banHER, next, is what they

call objective case–and that is all.”

�Der Werwolf“ – sprach der gute

Mann, �des Werwolfs, Genitiv sodann, dem Werwolf, Dativ, wie man’s

nennt, den Werwolf, - damit hat’s ein

End!“ “El hechiuno” le explica, “el hechidos” después indi-

ca, “el hechitres” dice al fin, “con eso basta ¡chiquitín!”

The banshee marveled at the cases and writhed with pleasure, making

faces, but said: “You did not add, so far, the plural to the singular!”

Dem Werwolf schmeichelten die

Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“ El hechicero, conmovido por lo que capta su oído, al maestro ruega, encanta-

do, que le adjunte un quebra-

do.

The teacher, though, admitted

then that this was not within his ken. “While bans are frequent”, he ad-

vised, “A she cannot be plurized.”

Der Dorfschullehrer aber mus-

ste gestehn, dass er von ihr nichts

wusste. Zwar Wölfe gab’s in großer

Schar, doch �Wer“ gab’s nur im Sinu-

lar.

Pero el profesor confiesa que nada sabe con certeza: “Hay muchos hechos. He-

chiceros únicamente hay enteros.”

…  …  …

2/2

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned crea-

ture, said humbly “Thanks” and left the

teacher.

Der Wolf erhob sich tränenblind

– er hatte ja doch Weib und Kind! Doch da er keine Gelehrter

eben, so schied er dankend und erge-

ben.

El brujo se volvió muy tri-

ste: “Mi cara mitad, ¿no la vi-

ste?” Ya que no era sabio con gratitud se despidió.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en hispana de Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

http://www.poezio.net/version?poem-id=825&version-id=1660,1652,1656